OpenStudy (anonymous):

f(x)=4x^2-x-3. If f(x)=-3, what is x? I'm stuck on this problem. What do I do after 0=4x^2-x

OpenStudy (whpalmer4):

\[f(x) = 4x^2-x-3\]\[f(x) = -3\]\[4x^2-x-3=-3\]Add 3 to both sides\[4x^2-x=0\]Can you factor the left side?

OpenStudy (anonymous):

Yeah, the x.

OpenStudy (anonymous):

Or -x?

OpenStudy (whpalmer4):

@mathslover I think they are trying to find the value of x that makes f(x) = -3, not evaluate f(-3)...

mathslover (mathslover):

Oops thanks for correction, I think i should delete my comment. Sorry @whpalmer4 for mistake and thanks a lot for correcting it.You may continue

OpenStudy (anonymous):

So do I factor out an x or -x?

OpenStudy (whpalmer4):

You could factor out either one, of course. You'll just have a different order to the stuff in the parentheses. \[x(4x-1) = 0\]or\[-x(1-4x)=0\]You should be able to convince yourself that they are equivalent.

OpenStudy (whpalmer4):

In either case, to have a product that equals 0, one or more of the things being multiplied must also equal 0. Set them each equal to 0 and solve for the values of \(x\) that make it true.

OpenStudy (anonymous):

Do I set it as x=0, 4x-1=0?